# Cooperative Decentralized Backdoor Attacks on Vertical Federated Learning

### v1

@@ I Introduction
In this paper, we consider an underexplored scenario where there are multiple adversaries, and these adversaries have the capability of colluding over a graph topology to execute a coordinated backdoor attack. Prominent examples can be found in defense settings. For instance, suppose attackers have taken control of a few scattered but networked military assets responsible for e.g., communicating front-line conditions, such as drones or tactical mobile devices [16]. Taking advantage of this, the adversaries could collude to analyze and transmit data across a covert adversary-formed network topology. By pooling their individual observations, they can potentially infer sensitive information about the battlefield–such as troop numbers, logistics capabilities, or defensive positions. Collusion allows the adversaries to expand their own individual insights, potentially making it easier for them to mislead the central decision-making system (e.g., at a command center) or manipulate outcomes via a backdoor attack. Still, the adversaries might be unable to engage in maximal information exchange (i.e., forming a fully connected graph) due to e.g., resource constraints, geographical distances, and channel conditions. We need to understand how adversarial collusion impacts the attack potency, and the role played by adversarial connectivity.
@@ I-A Related Work
Despite these recent efforts, a major limitation of the existing approaches is that they rely on information sent from the server to conduct the backdoor attack, in addition to using it for VFL participation. This dependency can enable the server to implement defense mechanisms, particularly during the label inference phase, which can significantly limit the effectiveness of the attack [35]. While research on bypassing the use of server-received information in backdooring VFL exists, it is limited to only binary classification tasks [36]. Additionally, the aforementioned studies [28, 29, 30, 31, 36] concentrate on the classic two-party VFL scenario with a single adversary, which fits them into the cross-silo FL context [9], leading to a limited understanding of VFL backdoors in networks where multiple adversaries may collude to carry out the attack. In particular, unlike cross-silo FL, which typically involves a few participants such as large organizations, cross-device FL often encompasses numerous distributed devices collaborating to construct a global model, thereby significantly increasing the potential for security malfunctions [1, 7]. These limitations lead us to investigate the following two research questions (RQs):
@@
Overview of approach. In this work, we develop a novel backdoor VFL strategy that addresses the above questions. To answer RQ1, we introduce a methodology for an adversary to locally infer and generate datapoints of the target label for attack. Our approach leverages Variational Autoencoders (VAE) [37] and triplet loss metric learning [38] to determine which samples should receive trigger embeddings, avoiding leveraging server gradients. To answer RQ2, we employ the graph topology of adversarial devices to conduct cooperative consensus on which samples should be implanted with triggers. In this regard, we show both empirically (Fig. 5(d) & Fig. 6) and theoretically (Sec. V) that the effectiveness of the attack and gradient perturbation is dependent on this graph topology. We also develop an intensity-based triggering scheme and two different methods for partitioning these triggers among adversaries, leading to a more powerful backdoor injection than existing attacks.
I-B Outline and Summary of Contributions
@@
  • We conduct extensive experiments comparing the performance of our attack against the state-of-the-art [28, 29] on four image classification datasets. Our results show that despite not using server information, we obtain a $30\%$ higher attack success rate for comparable main task performance. We also show an added advantage of our decentralized attack in terms of improved robustness to noising defenses at the server. We also demonstrate that higher adversarial graph connectivity yields improved attack success rate with our method, thus corroborating our theoretical claims (Sec. VI).
@@ III Attack Methodology
In order to execute the backdoor attack in VFL, adversaries need to (i) identify datapoints belonging to the target label (Sec. III-A) and (ii) implant triggers on the corresponding datapoints to induce misclassification (Sec. III-B). We present the methodology for these processes in this section, and give more specific algorithmic procedures for them in Sec. IV. Our label inference methodology is summarized in Fig. 2, whose individual phases we explain in greater detail in the following section (Sec. III-A).
@@ III-A1 Feature Sharing
As in existing work [29, 13], we assume that the adversaries possess labels for a small (e.g., $<1$%) set of the datapoints. Even so, when numerous clients each hold a small feature partition of the samples (Feature partition block of Fig. 2), extracting meaningful information without employing gradients from the server (which we aim to avoid, as discussed in Sec. I) becomes challenging. This difficulty arises due to the presence of irrelevant features within sample partitions, e.g., a blank background.
@@ III-A2 VAE and Metric-based Label Inference
[…]
@@
where $10<\lambda<$ 1 captures the importance weight of each individual term.
@@ III-B Trigger Embedding
In designing the trigger, ideally, we aim for the poisoned datapoints to produce embeddings that are as close as possible to embeddings of non-poisoned embeddings:
$\sigma_{m}^{\star}=\underset{\sigma_{m}}{\arg\min}\left\|\mathbb{E}[h_{m}^{\text{pois}}]-\mathbb{E}[h_{m}^{\text{target}}]\right\|_{2},$ (5)
where $\mathbb{E}[h_{m}^{\text{pois}}]$ is the expectation over feature embeddings produced from datapoints $i$ implanted with the trigger $\sigma_{m}$ and $\mathbb{E}[h_{m}^{\text{target}}]$ is the expectation over feature embeddings produced from clean datapoints $i^{\prime}$ belonging to the target label. If adversary $m$ had access to the server’s loss function, it may be possible to incorporate an approximation of (5) directly into the adversary’s local VFL update. However, we are considering a black-box scenario. Hence, to emulate the behavior of $\sigma_{m}^{\star}$, we develop an intensity-based triggering scheme that enhances the background value of the trigger by an adversary-defined intensity value $\gamma$, detailed in Sec. IV-B. In this way, the trigger becomes more prominent within the datapoint. Embedding the target datapoints with this intensified trigger, we can pull the backdoored ones closer to embeddings of the target label, making it harder for the server to distinguish between them.
@@ IV-A1 Training VAE $\phi_{m}^{\text{VAE}}$ and Auxiliary Classifier $\phi_{m,\mu}$
After training the VAE, $\phi_{m,\mu}$ take in the $\mu$-embeddings of the datapoints from $\mathcal{D}_{m}$ with unknown labels as input (Line 3 of Alg. 1), and populate $\mathcal{D}_{m}^{(p)}$, the set of locally inferred target datapoints. $\mathcal{D}_{m}^{(p)}$ is initially the same as $\widehat{\mathcal{D}}_{m}^{\text{target}}$ in terms of indices. The datapoints are added to the set if the maximum prediction probability corresponds to the target label $y_{t}$ and is greater than a confidence threshold $\beta$, i.e., $\max(\hat{y})\geq\beta$ and $y_{t}=\text{arg}\max_{c}(\hat{y}_{c})$ (Lines 5-7 of Alg. 1). In this way, only datapoints $\phi_{m,\mu}$ is confident about will be considered as targets.
@@ V Convergence Analysis
The loss function $F({\theta})$ in (1) is $L$-smooth, meaning that for any $x$ and $y$, we have $F(y)\leq F(x)+\langle\nabla F(x),y-x\rangle+\frac{L}{2}\|y-x\|^{2}$.
@@
The mini-batch gradient $\nabla_{\theta_{k}}\mathcal{L}$ is an unbiased estimate of $\nabla_{\theta_{k}}F({\theta})$, and $\mathbb{E}\|\nabla_{\theta_{k}}\mathcal{L}-\nabla_{\theta_{k}}F({\theta})\|^{2}\leq\Gamma,\forall k\in\mathcal{K}$, where $\Gamma$ is the variance of $\nabla_{\theta_{k}}\mathcal{L}$.
@@
There is an upper bound for the gradient perturbation from adversaries, i.e., $\mathbb{E}\|\nabla^{a}_{\theta_{k}}\mathcal{L}-\nabla_{\theta_{k}}\mathcal{L}\|^{2}\leq\delta(\rho),\forall k\in\mathcal{K}$, where $\nabla^{a}_{\theta_{k}}\mathcal{L}$ denotes the perturbed gradient and $\rho$ is a measure of connectivity for the graph $G$.
Assumptions 1 and 2 are common in literature [45, 46], while Assumption 3 is introduced to characterize the perturbation induced by the attack. Intuitively, we expect that $\delta(\rho)$ is an increasing function of $\rho$: a higher graph connectivity for collusion implies a higher probability that more datapoints will be targeted, leading to a larger perturbation. We will corroborate this relationship experimentally in Sec. VI, where $\rho$ is taken as the second smallest eigenvalue of the Laplacian matrix of $G$ (the algebraic connectivity or Fiedler value) [47], which often increases with average node degree.
@@ VI Numerical Experiments
In this section, we evaluate the performance of our proposed approach on various benchmark datasets. We compare our performance with two state-of-the-art backdoor VFL attack methods discussed in Sec. I-A: BadVFL [28] and VILLAIN [29].
@@ VI-A Simulation Setup
TABLE II: Network architecture and hyper-parameters.
@@
MNIST FMNIST CIFAR-10 SVHN
Margin $\kappa$ $0.4$ $0.25$ $0.2$ $0.2$
Poisoning budget $\zeta$ $1\%$ $1\%$ $1\%$ $1\%$
Intensity factor $\gamma$ $20$ $30$ $30$ $20$
Confidence threshold $\beta$ $0.999$ $0.999$ $0.9985$ $0.99995$
##\## Auxiliary datapoints $360$ $360$ $350$ $560$
[…]
VAE Latent Dimension $32$ $64$ $512$ $256$
We perform experiments on the MNIST [49], Fashion-MNIST (FMNIST) [50], CIFAR-10 [51], and Street View House Numbers (SVHN) [52] datasets. We consider fully-connected VAEs for MNIST and Fashion-MNIST, a Convolutional VAE (CVAE) with 4 layers each for the decoder and encoder for CIFAR-10, and a 4-layer encoder and 3-layer decoder CVAE for SVHN. For the bottom model, MNIST and Fashion-MNIST adopt a two layer Convolutional Neural Network (CNN), and CIFAR-10 has the same architecture as the encoder of the VAE. SVHN has the same bottom model as CIFAR-10. For the top model, we adopt a two-layer fully-connected network, which trains using the Adam optimizer. Further details are given in Table II.
In our experiments, unless stated otherwise, there are 10 clients with 5 adversaries, with the adversaries utilizing trigger Method 1 from Sec. IV-B by default. Moreover, since the baselines do not consider adversary graphs in their methodology, we by default consider a fully-connected graph for fair comparison. In addition, note that both baselines assume only one adversary, and we extend their method to multiple adversaries by adding majority voting and trigger splitting to their label inference and attack processes. All experiments were conducted on a server with a 40GB NVIDIA A100-PCIE GPU and 128GB RAM, using PyTorch for neural network design and training.
@@ VI-B Competitive Analysis with Baselines
Figure 4: Attack Success Rate (ASR) and Clean Data Accuracy (CDA) for MNIST, Fashion-MNIST, CIFAR-10, and SVHN. The proposed method converges to a higher ASR value than the baselines (BadVFL [28] and VILLAIN [29]) due to (1) having a higher label inference accuracy as seen in Fig. 5(a) and (2) having an intensity based trigger that makes it easier for the server to draw the association between the target label and trigger.
@@
Attack Success Rate (ASR) and Clean Data Acc. (CDA): First, we assess the accuracy of the backdoor attack and the main task and compare the performance of our approach to the baselines developed in [28] and [29]. We define the accuracy of misclassifying a poisoned datapoint as the target label as Attack Success Rate (ASR) and the accuracy of the regular main task as Clean Data Accuracy (CDA), which is the same notation utilized in [29]. We intend to show a high ASR, indicating a successful backdoor, while keeping the CDA close to the baselines, thereby showing the CDA is not significantly affected.
@@
Our results are given in Fig. 4. We first note the superiority of our approach compared to the baselines: our attack achieves higher ASR across all four datasets, while the CDA stays relatively constant with the CDAs of the baselines. There are three reasons for this. Firstly, due to better label inference performance (outlined in following section), the adversaries are more accurately poisoning datapoints belonging to the target, making it more likely for the server to draw an association between the target label and the trigger. Secondly, the intensity-based triggers are more easily captured by the bottom-models, making the server more reliant on it for classification. Lastly, the samples generated by the adversary VAEs follow the same general patterns and features of the target label, meaning that it is still learning the overall structure of the clean features properly when poisoned datapoints are present, keeping both the ASR and CDA high.
We note that our method requires a smaller trigger area compared to BadVFL [28] to successfully carry out the attack. Additionally, BadVFL is sensitive to its initial known datapoint, causing the backdoor to fail during some runs, accounting for the high variation in the averages. This can be seen in Table III, where the ASR average is low for MNIST, CIFAR-10, and SVHN due to a poor initial known datapoint being used often. On the other hand, VILLAIN [29] must wait for a period of time before label inference and attacking can begin, since the method requires a well-trained bottom and top model to carry out an attack. Lastly, the baselines, particularly VILLAIN, slightly suffer from catastrophic forgetting, where the backdoor task is slowly forgotten over continuous iterations [53], whereas, our proposed attack consistently performs well.
Label Inference: We compare the accuracy of our label inference to the baselines [28, 29] in Fig. 5(a). We see that the proposed method of utilizing a combination of metric learning and VAEs for label inference outperforms the baselines in terms of label inference accuracy for all datasets, while not relying on the server information. Note that the proposed attack can still be quite successful (Fig. 4) even when the label inference accuracy is not very high (e.g., with the CIFAR-10 and SVHN datasets). This indicates that some margin of error in label inference can be acceptable, provided that the target data points significantly outnumber the combined total of other labels.

### v2

@@ I Introduction
In this paper, we consider an underexplored scenario where there are multiple adversaries, and these adversaries have the capability of colluding over a graph topology to execute a coordinated backdoor attack. In this setting, the adversaries gain control of a set of client nodes and establish an ad-hoc local network among themselves for cooperation. Prominent examples can be found in defense settings. For instance, in a contested region, attackers could take control of a few scattered military assets responsible for e.g., communicating front-line conditions, such as drones or tactical mobile devices [16]. Such control could be achieved through a variety of means, for example, exploiting vulnerabilities over control links to hijack drones [17] or injecting malicious malware onto the devices [18]. Taking advantage of this, the adversaries could begin colluding to analyze and transmit data across a local adversary-formed network topology. To form this graph, the adversaries could employ device-to-device (D2D) communication protocols, which have proliferated in 5G wireless and have shown benefit in distributed learning, allowing devices to synchronize model parameters and/or gain a better estimate of the overall data distribution [19, 20, 21]. This decentralized topology allows for the exchange of information between adversaries that might not be readily available at an individual level, i.e., a more complete view of each sample. By pooling their individual observations, they can potentially infer sensitive information about the battlefield–such as troop numbers, logistics capabilities, or defensive positions. Moreover, this making it easier for them to mislead the central decision-making system via a backdoor attack (e.g., inducing the command center to misclassify an enemy fighter as a benign node).
Still, the adversaries might be unable to engage in maximal information exchange (i.e., forming a full mesh graph) due to e.g., resource constraints, geographical distances, and channel conditions preventing formation of certain D2D links. We thus need to understand how adversarial collusion impacts the attack potency, and the role played by adversarial connectivity. Moreover, following the common assumption in the literature that the adversaries have full control over the compromised node(s) [22], it is important to consider the full range of adversarial capabilities. By doing so, we do not underestimate the adversary’s capabilities and insights, which could otherwise lead to overconfidence in the system’s security.
@@ I-A Related Work
Despite these recent efforts, a major limitation of the existing approaches is that they rely on information sent from the server to conduct the backdoor attack, in addition to using it for VFL participation. This dependency can enable the server to implement defense mechanisms, particularly during the label inference phase, which can significantly limit the effectiveness of the attack [41]. While research on bypassing the use of server-received information in backdooring VFL exists, it is limited to only binary classification tasks [42]. Additionally, the aforementioned studies [34, 35, 36, 37, 42] concentrate on the classic two-party VFL scenario with a single adversary, which fits them into the cross-silo FL context [9], leading to a limited understanding of VFL backdoors in networks where multiple adversaries may collude to carry out the attack. In particular, unlike cross-silo FL, which typically involves a few participants such as large organizations, cross-device FL often encompasses numerous distributed devices collaborating to construct a global model, thereby significantly increasing the potential for security malfunctions [1, 7].
I-B Research Questions and Approach Overview
These limitations lead us to investigate the following two research questions (RQs) in this work:
@@ I-B Research Questions and Approach Overview
Overview of approach. We develop a novel backdoor VFL strategy that addresses the above questions. To answer RQ1, we introduce a methodology for an adversary to locally infer and generate datapoints of the target label for attack. Our approach leverages Variational Autoencoders (VAE) [43] and triplet loss metric learning [44] to determine which samples should receive trigger embeddings, avoiding leveraging server gradients. To answer RQ2, we employ the graph topology of adversarial devices to conduct cooperative consensus on which samples should be implanted with triggers. In this regard, we show both empirically (Fig. 5(d) & Fig. 6) and theoretically (Sec. V) that the effectiveness of the attack and gradient perturbation is dependent on this graph topology. We also develop an intensity-based triggering scheme and two different methods for partitioning these triggers among adversaries, leading to a more powerful backdoor injection than existing attacks.
I-C Outline and Summary of Contributions
@@
  • We conduct extensive experiments comparing the performance of our attack against the state-of-the-art [34, 35] on five image classification datasets. Our results show that despite not using server information, we obtain a $30\%$ higher attack success rate for comparable main task performance. We also show an added advantage of our decentralized attack in terms of improved robustness to noising defenses at the server. We also demonstrate that higher adversarial graph connectivity yields improved attack success rate with our method, thus corroborating our theoretical claims (Sec. VI).
@@ III Attack Methodology
To execute the backdoor attack in VFL, adversaries need to (i) identify datapoints belonging to the target label (Sec. III-A) and (ii) implant triggers on the corresponding datapoints to induce misclassification (Sec. III-B). We present the methodology for these processes in this section, and give more specific algorithmic procedures in Sec. IV.
@@ III-A1 Feature Sharing
As in existing work [35, 13], we assume that the adversaries possess labels for a small (e.g., <1<1%) set of the datapoints. Even so, when numerous clients each hold a small feature partition of the samples (Feature partition block of Fig. 2), extracting meaningful information without employing gradients from the server (which we aim to avoid, as discussed in Sec. I) becomes challenging. This difficulty arises due to the presence of irrelevant features within sample partitions, e.g., a blank background.
@@ III-A2 VAE and Metric-based Label Inference
However, $p(X_{m}^{(i)})=\int p_{m,d}(X_{m}^{(i)}|z_{m}^{(i)})p(z_{m}^{(i)})dz_{m}^{(i)}$ is computationally intractable, making it unrealistic to calculate the term directly. Therefore, rather than maximizing $p_{m,d}$ directly, the VAE employs its encoder $q_{m,e}$ as an approximate model which outputs a mean $\mu$ and standard deviation $\sigma$, reducing the latent space to a univariate Gaussian $\mathcal{N}(\mu,\sigma^{2})$. The error can be captured in a KL divergence-based objective [46] which measures the difference between two probability distributions, $q_{m,e}(z_{m}^{(i)})$ and $p(z_{m}^{(i)})$, denoted as $D_{\text{KL}}(q_{m,e}(z_{m}^{(i)})\|p(z_{m}^{(i)}))$. This term, when included in the loss function, can encourage the latent space to be closer to a standard normal distribution, allowing for random sampling from the latent space for datapoint generation.
@@
where $0<\lambda<1$ captures the importance weight of each individual term.
@@ III-A3 Collaborative Inference
Each of these label inference processes, e.g., feature sharing, VAE training/inference, and collaborative inference, take place before the VFL training process begins. This emphasizes one of our contributions: developing a backdoor attack where gradients from the server are not utilized for label inference. As, we will see in Sec. VI-B, this enhances robustness against server-side gradient noise injection defenses. Further, by minimizing the degree to which adversary behavior will deviate from the VFL training protocol, the possibility of the system detecting the attack (through e.g., anomalous communication, computation, or energy consumption) becomes small.
@@ III-B Trigger Embedding
In designing the trigger, we aim for the poisoned datapoints to produce embeddings that are as close as possible to embeddings of non-poisoned embeddings. This can be thought of mathematically as looking for triggers that will minimize $\|\mathbb{E}[h_{m}^{\text{pois}}]-\mathbb{E}[h_{m}^{\text{target}}]\|_{2}$, where $\mathbb{E}[h_{m}^{\text{pois}}]$ is the expectation over feature embeddings produced from datapoints $i$ implanted with the trigger and $\mathbb{E}[h_{m}^{\text{target}}]$ is the expectation over feature embeddings produced from clean datapoints $i^{\prime}$ belonging to the target label. If adversary $m$ had access to the server’s loss function, it may be possible to incorporate (an approximation of) this norm difference directly into the adversary’s local VFL update. However, we are considering a black-box scenario. Hence, to emulate the desired trigger behavior, we complement the data substitution and selective poisoning process with an intensity-based triggering scheme. Detailed in Sec. IV-B, this scheme enhances the background value of the trigger by an adversary-defined intensity value $\gamma$, so that the trigger becomes more prominent within the datapoint. Thus, on the one hand, $\widetilde{x}_{m}^{(i)}$ presents the server with more variations of the target label, causing the server to rely more heavily on the consistent trigger pattern. Then, combining these harder-to-identify background features with the $\gamma$-enhanced trigger induces heavier reliance of the server on the trigger while minimizing main task performance degradation (when the trigger is not present).
The VFL system relies on the bottom models’ outputted latent embeddings for classification, with updates to these local models aiming to optimize performance. By embedding the target datapoints with an intensified trigger, we can pull the embeddings $h_{m}^{\text{pois}}$ of the backdoored samples closer to embeddings $h_{m}^{\text{target}}$ of the target label, making it harder for the server to distinguish between them.
@@ IV-A1 Training VAE $\phi_{m}^{\text{VAE}}$ and Auxiliary Classifier $\phi_{m,\mu}$
After training the VAE, $\phi_{m,\mu}$ takes in the $\mu$-embeddings of the datapoints from $\mathcal{D}_{m}$ with unknown labels as input (Line 3 of Alg. 1), and populate $\mathcal{D}_{m}^{(p)}$, the set of locally inferred target datapoints. $\mathcal{D}_{m}^{(p)}$ is initially the same as $\widehat{\mathcal{D}}_{m}^{\text{target}}$ in terms of indices. The datapoints are added to the set if the maximum prediction probability corresponds to the target label $y_{t}$ and is greater than a confidence threshold $\beta$, i.e., $\max(\hat{y})\geq\beta$ and $y_{t}=\text{arg}\max_{c}(\hat{y}_{c})$ (Lines 5-7 of Alg. 1). In this way, only datapoints $\phi_{m,\mu}$ is confident about will be considered as targets.
@@ V Convergence Analysis
The loss function $F({\theta})$ in (1) is $L$-smooth, meaning that for any $x$ and $y$, we have
$F(y)\leq F(x)+\langle\nabla F(x),y-x\rangle+\frac{L}{2}\|y-x\|^{2}.$
@@
The mini-batch gradient $\nabla_{\theta_{k}}\mathcal{L}$ is an unbiased estimate of $\nabla_{\theta_{k}}F({\theta})$, and
$\mathbb{E}\|\nabla_{\theta_{k}}\mathcal{L}-\nabla_{\theta_{k}}F({\theta})\|^{2}\leq\Gamma,\forall k\in\mathcal{K},$
where $\Gamma$ is the variance of $\nabla_{\theta_{k}}\mathcal{L}$.
@@
There is an upper bound for the gradient perturbation from adversaries, i.e.,
$\mathbb{E}\|\nabla^{a}_{\theta_{k}}\mathcal{L}-\nabla_{\theta_{k}}\mathcal{L}\|^{2}\leq\delta(\rho),\forall k\in\mathcal{K},$
where $\nabla^{a}_{\theta_{k}}\mathcal{L}$ denotes the perturbed gradient and $\rho$ is a measure of connectivity for the graph $G$.
[…]
@@ VI Numerical Experiments
In this section, we evaluate the performance of our proposed approach on various datasets. We compare our performance with two state-of-the-art backdoor VFL attack methods discussed in Sec. I-A: BadVFL [34] and VILLAIN [35].
@@ VI-A Simulation Setup
TABLE II: Network architecture and hyperparameters for each dataset.
@@
MNIST FMNIST
Margin $\kappa$ $0.4$ $0.25$
Poisoning budget $\zeta$ $\% 11\%$
Intensity factor $\gamma$ 2020 $30$
Confidence threshold $\beta$ $0.999$ $0.999$
#\# Auxiliary samples $360$ $360$
$\%$ Auxiliary as Target $\% 1616\%$
VAE Latent Dimension $32$ $64$
Parameters Dataset
CIFAR-10 SVHN CIFAR100-20
Margin $\kappa$ $0.2$ $0.2$ $0.2$
Poisoning budget $\zeta$ $\% 11\%$ $0.5\%$
Intensity factor $\gamma$ $30$ 2020 $95$
Confidence threshold $\beta$ $0.9985$ $0.99995$ $0.99999$
#\# Auxiliary samples $350$ $560$ $487$
$\%$ Auxiliary as Target $14\%$ $12.5\%$ $15.4\%$
VAE Latent Dimension $512$ $256$ $512$
We perform experiments on the MNIST [55], Fashion-MNIST (FMNIST) [56], CIFAR-10 [57], Street View House Numbers (SVHN) [58], and CIFAR-100 (CIFAR100-20, coarse label version) [57, 59, 60] datasets. We consider fully-connected VAEs for MNIST and FMNIST, a Convolutional VAE (CVAE) with 4 layers each for the decoder and encoder for CIFAR-10 and CIFAR100-20, and a 4-layer encoder and 3-layer decoder CVAE for SVHN. For the bottom model, MNIST and FMNIST adopt a two layer Convolutional Neural Network (CNN), with CIFAR-10 and CIFAR100-20 having the same architecture as the encoder of the VAE. SVHN has the same bottom model as CIFAR-10. For the top model, we adopt a two-layer fully-connected network, which trains using the Adam optimizer [61]. Further details are given in Table II.
In our experiments, unless stated otherwise, there are 10 clients with 5 adversaries, with the adversaries utilizing trigger Method 1 from Sec. IV-B by default. Moreover, since the baselines do not consider adversary graphs in their methodology, we by default consider a fully-connected graph for fair comparison. In addition, note that both baselines assume only one adversary, and we extend their method to multiple adversaries by adding majority voting and trigger splitting to their label inference and attack processes. All experiments were conducted on a server with a 40GB NVIDIA A100-PCIE GPU and 128GB RAM, using PyTorch [62] for neural network design and training.
@@ VI-B Competitive Analysis with Baselines
Figure 4: Attack Success Rate (ASR) and Clean Data Accuracy (CDA) for MNIST, Fashion-MNIST, CIFAR-10, SVHN, and CIFAR100-20. The proposed method converges to a higher ASR value than the baselines (BadVFL [34] and VILLAIN [35]) due to (1) having a higher label inference accuracy as seen in Fig. 5(a) and (2) having an intensity based trigger that makes it easier for the server to draw the association between the target label and trigger.
@@
Attack Success Rate (ASR) and Clean Data Acc. (CDA): First, we assess the accuracy of the backdoor attack and the main task and compare the performance of our approach to the BadVFL and VILLAN baselines. We define the accuracy of misclassifying a poisoned datapoint as the target label as Attack Success Rate (ASR) and the accuracy of the regular main task as Clean Data Accuracy (CDA), which is the same notation utilized in [35]. We intend to show a high ASR, indicating a successful backdoor, while keeping the CDA close to the baselines, indicating the main task is not significantly affected.
@@
Our results are given in Fig. 4. We note the superiority of our approach compared to the baselines: our attack achieves higher ASR across all settings, while the CDA stays relatively constant with the CDAs of the baselines. There are three reasons for this. First, due to better label inference performance (outlined in following section), the adversaries are more accurately poisoning datapoints belonging to the target, making it more likely for the server to draw an association between the target label and the trigger. Secondly, the intensity-based triggers are more easily captured by the bottom-models, making the server more reliant on it for classification. Third, the samples generated by the adversary VAEs follow the same general patterns and features of the target label, meaning that it is still learning the overall structure of the clean features properly when poisoned datapoints are present, keeping both the ASR and CDA high. Note that our findings are consistent on each of the datasets; compared to MNIST, Fashion-MNIST, CIFAR-10, and SVHN, the CIFAR100-20 experiments validate our results on a larger and more complex dataset. This corroborates the efficacy of our proposed method in implanting a successful backdoor on a wide range of learning scenarios.
We also note that our method requires a smaller trigger area compared to BadVFL [34] to successfully carry out the attack. Additionally, BadVFL is sensitive to its initial known datapoint, causing the backdoor to fail during some runs, accounting for the high variation in the averages. As seen in Fig. 4, the ASR average is low for CIFAR100-20 and SVHN due to a poor initial known datapoint being used often. On the other hand, VILLAIN [35] must wait for a period of time before label inference and attacking can begin, since the method requires a well-trained bottom and top model to carry out an attack. Lastly, the baselines, particularly VILLAIN, slightly suffer from catastrophic forgetting, where the backdoor task is slowly forgotten over continuous iterations [63], whereas, our proposed attack consistently performs well.
Label Inference: We compare the accuracy of our label inference to the baselines [34, 35] in Fig. 5(a). We see that the proposed method of utilizing a combination of metric learning and VAEs for label inference has the best performance in terms of label inference accuracy across the datasets, while not relying on the server information. Note that the proposed attack can still be quite successful (Fig. 4) even when the label inference accuracy is not very high (e.g., with the CIFAR-10 and SVHN datasets). This indicates that some margin of error in label inference can be acceptable, provided that the target data points significantly outnumber the combined total of other labels. As in Fig. 4, the inclusion of these five datasets validates the label inference accuracy module of our methodology across a wide range of learning scenarios. Specifically, while our proposed method achieves high label inference accuracy on benchmark datasets such as CIFAR-10 and MNIST, the results on the CIFAR100-20 dataset further demonstrate its robustness and scalability on more complex and diverse settings.